OpenStudy (anonymous):

Given that the P(A) = 0.32, P(B) = 0.34, and P(AUB)= 0.56, find the P( B ' | A' ).

OpenStudy (anonymous):

p(A' intersection B')=p(A)+p(B)-p(A intersection B) p(A intersection B) u can get from the equations

OpenStudy (anonymous):

P( B ' | A' ).=P(A' intersection B')/P(B')

OpenStudy (anonymous):

can you solve it now

OpenStudy (zarkon):

\[P(A'\cap B')=1-P(A\cup B)\] \[P(B'|A')=\frac{P(B'\cap A')}{P(A')}\]

OpenStudy (zarkon):

thus \[P(B'|A')=\frac{P(B'\cap A')}{P(A')}=\frac{1-P(A\cup B)}{1-P(A)}\]